# Mechanical hysterons with tunable interactions of general sign

### v2

@@ Appendix A Appendix
Spring stiffnesses.— We measure the spring stiffnesses from force-versus-displacement data that we obtain by anchoring the springs to a known mass on an electronic balance, and recording the force on the balance as the top of the spring is extended and then retracted (Fig. 1B). These measurements show that the springs behave linearly over the range of extensions used in the experiments.
@@
Torques for general coupling.— Here we derive expressions for the torques in the small-$\theta$ model, in the general case where the two springs connecting hysterons $i$ and $j$ are not necessarily paired with equal spring constants or rest lengths, and where they may not have symmetrical mounting positions. We begin by considering a single coupling spring of stiffness $k$ and rest length $x_{0}$ that mounts to rotors $1$ and $2$ at positions $\ell_{1}$ and $\ell_{2}$ from the fulcrums of the rotors, respectively, where the signs of $\ell_{1}$, $\ell_{2}$ indicate the side of the fulcrum where they attach. In the limit of small $\theta_{1}$, $\theta_{2}$ this coupling spring exerts a torque on hysteron 1 equal to:
$\tau=k(\ell_{1}w+\ell_{1}\ell_{2}(\theta_{2}-\theta_{1}))\left(1-\frac{x_{0}}{L_{0}}\right)\\ +\frac{k\ell_{1}w^{2}x_{0}}{L_{0}^{3}}(\ell_{2}\theta_{2}-\ell_{1}\theta_{1})\ ,$ (7)
@@
where $L_{0}=\sqrt{(\ell_{2}-\ell_{1})^{2}+w^{2}}$ is the length of the spring when $\theta_{1}=\theta_{2}=0$. With this result, it is straightforward to write down the net torque $\tau_{12}=\tau+\tau^{\prime}$ on rotor 1 for the general case where a second spring of stiffness $k^{\prime}$ and rest length $x_{0}^{\prime}$ is mounted at distinct positions $\ell_{1}^{\prime}$ and $\ell_{2}^{\prime}$. The expression will in general depend on the properties and mounting positions of both springs, and will contain terms proportional to $\ell_{1}$, $\ell_{1}^{2}$, and $\ell_{1}\ell_{2}$, for both unprimed and primed $\ell$.
Pairing the springs as in Fig. 1D with $k=k^{\prime}$ and $x_{0}=x_{0}^{\prime}$ greatly simplifies the interaction. Mounting the springs in a “parallel” configuration, $\ell=\ell_{1}=\ell_{2}=-\ell_{1}^{\prime}=-\ell_{2}^{\prime}$, we obtain:
$\tau_{12}=2k\ell^{2}(\theta_{2}-\theta_{1})\ ,$ (8)
where the dependence on $x_{0}=x_{0}^{\prime}$ has vanished, and the dependence is on $\ell^{2}$ alone.
@@
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})\\ -2k\ell^{2}\frac{x_{0}}{\sqrt{4\ell^{2}+w^{2}}}\left(2\theta_{1}-\frac{4\ell^{2}}{4\ell^{2}+w^{2}}(\theta_{2}+\theta_{1})\right)\ .$ (9)
Specializing to $x_{0}\ll w$ simplifies this expression to:
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})\ ,$ (10)
which is equal and opposite to Eq. 8.

### v3

@@ Abstract
Spring stiffnesses. We measure the spring stiffnesses from force-versus-displacement data that we obtain by anchoring the springs to a known mass on an electronic balance, and recording the force on the balance as the top of the spring is extended and then retracted (Fig. 1B). These measurements show that the springs behave linearly over the range of extensions used in the experiments.
Abstract hysteron simulations. We verified the period-2 behavior of Fig. 5A using the open-source hysteron software package Keim and Paulsen (2020).
Acknowledgements. I thank Nathan Keim for suggesting how to apply a global drive, for providing the parameters used in Fig. 5A, and for many fruitful discussions. I am grateful to Zachariah Schrecengost for identifying how the interactions could be treated at arbitrary rotor angles. I thank Vidyesh Anisetti, Eadin Block, Christian Santangelo, and Jennifer Schwarz for contributing to early hysteron designs, and I thank Martin van Hecke, Nathan Keim, Chloe Lindeman, Gentian Muhaxheri, and Dor Shohat for comments on the manuscript. This work was supported by Syracuse University research subsidy funds and was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
Supplementary Information for
“Mechanical hysterons with tunable interactions of general sign”
Joseph D. Paulsen
I Targeting specific transition graphs
We start by demonstrating four distinct transition graphs in a sequence where we alter only the mounting position of driving spring 2, thereby changing only $x_{2}$. Supplementary Movie 2 shows this sequence of behaviors, which correspond to transition graphs $i$-iviv in Fig. S1A. The experimental parameters are tabulated in Fig. S1C.
To understand how the system moves between these four transition graphs, we plug these parameters into the kinematic model, calculate the switching thresholds, and plot them in Fig. S1B. The switching thresholds of hysteron 1 are constant throughout this sequence, as they depend on the state of hysteron 2 but not its driving. The thresholds for hysteron 2 move in concert without changing their spacing, as changing $x_{2}$ shifts each threshold by $x_{2}$. Careful inspection of the switching thresholds in Fig. S1B shows that they indeed produce graphs $i$-iviv in Fig. S1A, faithful to the experiment.
Next, we access transition graph $v$ by decreasing $\Delta\theta_{1}$, increasing $\Delta\theta_{2}$, and changing $x_{2}$ (Supplementary Movie 3). We access transition graph $vi$ by setting $\Delta\theta_{1}$ = $\Delta\theta_{2}$ and changing $x_{2}$ once again (Supplementary Movie 4). This exhausts all possible transition graphs for two hysterons with cooperative interactions van Hecke (2021).
Three additional transition graphs can be obtained by using frustrated interactions van Hecke (2021); they are graphs viivii-ixix in Fig. S1A. Supplementary Movie 5 shows how we can realize these graphs in the experiment while varying only $x_{2}$ to navigate between them. All three graphs feature at least one “horizontal avalanches”. To understand how they occur, start in the -- state in graph viivii. The corresponding thresholds in Fig. S1B show that upon increasing $\gamma$, hysteron 2 switches first, as $\gamma_{2}^{+}(-)<\gamma_{1}^{+}(-)$. Now in the +-+ state, the next event is at $\gamma=\gamma_{1}^{+}(+)$ into ++++. But this state is not stable, as this $\gamma$ is below $\gamma_{2}^{-}(+)$, which brings the system immediately to ++-. An analogous process underlies each of the $\leftrightarrow+-$ avalanches in graphs viivii-ixix.
Among these graphs, perhaps the most exotic is $ix$, which navigates the sequence: --, +-+, ++-, ++++ under unidirectional driving. This behavior constitutes a 2-bit analog-to-digital converter van Hecke (2021). We now describe the process used to target this transition graph, to illustrate how behaviors may be rationally designed in our system. First, we need a strong coupling between the global drive and hysteron 1, so that it is ()(-) at small $\gamma$ and $(+)$ at large $\gamma$. We also need a strong frustrated interaction between hysterons 1 and 2 to navigate the $\leftrightarrow+-$ avalanches. In particular, the interaction of hysteron 1 on hysteron 2 must open up a central interval, $\gamma_{2}^{+}(-)<\gamma<\gamma_{2}^{-}(+)$ where the state of hysteron 2 depends only on the current state of hysteron 1. Figure S1D illustrates this condition, and an examination of graphs viivii-ixix in Fig. S1 shows that they satisfy it. We target this condition by selecting a large $k_{12}$ and a sufficiently large $\theta_{1}^{\pm}$; we find that the same $\pm 38.8^{\circ}$ used in the other designs works here too. Finally, a weak coupling of hysteron 2 to the global drive enables a transition into the ++++ state at large $\gamma$ and into -- at small $\gamma$. These considerations lead to a tentative design for the spring constants: $k_{1}\approx k_{12}\gg k_{2}$. Here we use $k_{1}=k_{12}=0.46$ N/cm and $k_{2}=0.07$ N/cm, which we found to suffice.
Two additional transition graphs are possible for “mixed” interactions, where the signs of $J_{ij}$ and $J_{ji}$ differ van Hecke (2021). We do not see a way to realize such interactions in our system.
II Torque from the driving rod
Here we derive the torque exerted by the driving rod on a hysteron. A driving spring of stiffness $k_{i}$ and rest length $x_{0}$ mounts to rotor $i$ at a position $L>0$ from the fulcrum of the rotor, with coordinates $\vec{r}_{L}=\{L\sin\theta_{i},L\cos\theta_{i}\}$. The other end of the driving spring connects to the driving rod at coordinates $\vec{r}_{\text{bar}}=\{\gamma+x_{i},-y\}$. The springs in the experiment obey Hooke’s Law: $F=k_{i}(x-x_{0})$, where $x=|\vec{r}_{\text{bar}}-\vec{r}_{L}|$. In vector form,
$\vec{F}=k(\vec{r}_{\text{bar}}-\vec{r}_{L})\left(1-\frac{x_{0}}{|\vec{r}_{\text{bar}}-\vec{r}_{L}|}\right)\ .$ (S1)
The torque on rotor $i$ is $\tau_{i}=-\vec{r}_{L}\times\vec{F}$, where use a minus sign when defining the torque, so that clockwise rotation of the rotors corresponds to increasing $\theta$. Computing the cross product by components, we obtain the exact expression:
$\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})\left(1-\frac{x_{0}}{\sqrt{(\gamma+x_{i}-L\sin\theta_{i})^{2}+(y+L\cos\theta_{i})^{2}}}\right)\ .$ (S2)
We use Eq. S2 when computing switching thresholds in the kinematic model.
When the rest length of the driving spring, $x_{0}$, is small compared to $L+y$, the torque is approximately:
$\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})\ .$ (S3)
We use Eq. S3 to derive Eqs. 3 and 4 in the main text and for the small-$\theta$ model.
III Torques between coupled hysterons
Here we derive expressions for the torque exerted by one hysteron on another. We begin with the most general case: arbitrary angles for the rotors, finite rest lengths for the springs, and arbitrary mounting positions of the springs on the rotors (Eq. S5). We then expand this result to leading order in $\theta_{1},\theta_{2}$ (Eq. S6). We then specialize this small-angle result to the “paired springs” that are used exclusively in the main text. The result for the “parallel” configuration turns out to be independent of the rest length of the springs (Eq. S7), whereas the “crossed” configuration requires one more limit to reach a similarly simple expression (Eq. S9).
III.1 Single coupling spring
@@
We consider a single coupling spring of stiffness $k$ and rest length $x_{0}$ that mounts to rotors $1$ and $2$ at positions $\ell_{1}$ and $\ell_{2}$ from the fulcrums of the rotors, respectively, where the signs of $\ell_{1}$, $\ell_{2}$ indicate the side of the fulcrum where they attach. The coordinates of the mounting position on rotor $1$ are $\vec{r}_{1}=\{\ell_{1}\sin\theta_{1},\ell_{1}\cos\theta_{1}\}$, and on rotor $2$ they are $\vec{r}_{2}=\{w+\ell_{2}\sin\theta_{2},\ell_{2}\cos\theta_{2}\}$. The springs in the experiment obey Hooke’s Law: $F=k(x-x_{0})$, where $x=|\vec{r}_{2}-\vec{r}_{1}|$. In vector form,
$\vec{F}=k(\vec{r}_{2}-\vec{r}_{1})\left(1-\frac{x_{0}}{|\vec{r}_{2}-\vec{r}_{1}|}\right)\ .$ (S4)
The torque on rotor $1$ is $\tau=-\vec{r}_{1}\times\vec{F}$, where use a minus sign when defining the torque, so that clockwise rotation of the rotors corresponds to increasing $\theta$. Computing the cross product by components, we obtain the exact expression:
$\tau=k\ell_{1}(w\cos\theta_{1}+\ell_{2}\sin(\theta_{2}-\theta_{1}))\left(1-\frac{x_{0}}{\sqrt{(w+\ell_{2}\sin\theta_{2}-\ell_{1}\sin\theta_{1})^{2}+(\ell_{2}\cos\theta_{2}-\ell_{1}\cos\theta_{1})^{2}}}\right)\ .$ (S5)
Expanding Eq. S5 in $\theta_{1}$ and $\theta_{2}$, we obtain the torque on hysteron 1 to leading order in $\theta_{1},\theta_{2}$:
$\tau=k\ell_{1}(w+\ell_{2}(\theta_{2}-\theta_{1}))\left(1-\frac{x_{0}}{L_{0}}\right)+k\ell_{1}(\ell_{2}\theta_{2}-\ell_{1}\theta_{1})\frac{w^{2}x_{0}}{L_{0}^{3}}+$...\ , (S6)
where $L_{0}=\sqrt{(\ell_{2}-\ell_{1})^{2}+w^{2}}$ is the distance spanned by the spring when $\theta_{1}=\theta_{2}=0$.
Equations S5 and S6 apply when rotor $2$ is to the right of rotor $1$. If rotor $2$ is on the left, sending $\ell_{1}\rightarrow-\ell_{1}$ and $\ell_{2}\rightarrow-\ell_{2}$ gives the relevant formulas, as this represents a $180^{\circ}$ rotation about the center of rotor $1$, which leaves both rotor angles and the coupling spring length unchanged.
III.2 Paired springs at small rotor angles
@@
Using Eq. S6, it is straightforward to write down the net torque $\tau_{12}=\tau+\tau^{\prime}$ on rotor 1 for the general case where a second spring of stiffness $k^{\prime}$ and rest length $x_{0}^{\prime}$ is mounted at distinct positions $\ell_{1}^{\prime}$ and $\ell_{2}^{\prime}$. Pairing the springs as in Fig. 1D with spring constants $k=k^{\prime}$ and $x_{0}=x_{0}^{\prime}$ greatly simplifies the form of the interaction. Mounting the springs in a “parallel” configuration, $\ell=\ell_{1}=\ell_{2}=-\ell_{1}^{\prime}=-\ell_{2}^{\prime}$, we find:
$\tau_{12}=2k\ell^{2}(\theta_{2}-\theta_{1})\ .$ (S7)
Notably, the net torque does not depend on the finite rest length of the springs, $x_{0}=x_{0}^{\prime}$, or the distance between rotors, $w$, due to the symmetry of such paired coupling springs.
@@
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})-2k\ell^{2}\frac{x_{0}}{\sqrt{4\ell^{2}+w^{2}}}\left(2\theta_{1}-\frac{4\ell^{2}}{4\ell^{2}+w^{2}}(\theta_{2}+\theta_{1})\right)\ .$ (S8)
Specializing to either $x_{0}\ll w$ or $x_{0}\ll\ell$ simplifies this expression to:
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})\ ,$ (S9)
which is equal and opposite to Eq. S7.
IV Targeting a latching behavior
In Fig. 3 of the main text we demonstrate a “latching” behavior, which had been introduced in the context of sheared amorphous solids in recent work by Lindeman et al. Lindeman et al. (2025). Here we describe how to rationally target this behavior using our mechanical hysterons. Lindeman et al. Lindeman et al. (2025) showed how latching is tied to the ordering of the switching thresholds. They found that latching requires:
$\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)<\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)\ ,$ (S10a)
$\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)\ .$ (S10b)
We now show how we translate this ordering into a concrete design in our mechanical system. We start by noting that Eq. S10a implies: $0<\gamma_{2}^{-}(+)-\gamma_{2}^{-}(-)<\gamma_{1}^{-}(+)-\gamma_{1}^{-}(-)$, which we rewrite as: $0<-J_{21}<-J_{12}$. We obtain this non-reciprocal frustrated interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them together with crossed springs following Eq. 6. Next, we note that adjusting $x_{2}$ moves all thresholds for hysteron 2 in concert without altering their spacing. We thus move $x_{2}$ until $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$ and $\gamma_{1}^{-}(-)<\gamma_{2}^{-}(-)$. Then, we fulfill $\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)$ by moving $x_{1}$ and $x_{2}$ together by equal amounts until those thresholds straddle $\gamma=0$. Finally, $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$ turns out to be satisfied already through the larger angular interval of hysteron 2.
V Counting driving cycles
Here we provide additional details regarding the accumulator design in the main text (Fig. 4). The basic idea of our design can be verified on a computer Keim and Paulsen (2020) by applying cyclic driving of any fixed amplitude $\gamma_{0}$, setting thresholds $\gamma_{2m-1}^{+}>-\gamma_{2m-1}^{-}=\gamma_{0}/2$ for the odd hysterons and $\gamma^{\pm}_{2m}=-\gamma^{\mp}_{2m-1}$ for the even hysterons (where $m=1,2,3$...), and setting nearest-neighbor interactions of strength $J_{i,i+1}=-\gamma_{0}$. The thresholds for hysteron 1 should be raised by $\gamma_{0}$ to allow a domain wall to nucleate there.
More precisely, our accumulator design detects cycles of intermediate amplitude. Small-amplitude driving $\gamma(t)$ that is bounded by $\max(\gamma_{i}^{-})<\gamma(t)<\min(\gamma_{i}^{+})$ will not change the state of any hysteron, whereas large-amplitude driving that overpowers the frustrated interactions in the bulk can send the system to its final absorbing state. Intermediate-amplitude cycles advance the domain wall by two hysterons (one per half-cycle).
Figure S1: Suite of behaviors for an interacting hysteron pair. A. Transition graphs obtained in the experiment, showing all states and transitions accessible from the initial state, -- (Supplementary Movies 2-5). Graphs $i$ and iiii are “Preisach” graphs that do not require interactions van Hecke (2021), whereas graphs iiiiii-vivi require cooperative interactions. To show how one can move between different transition graphs, we demonstrate them all using cooperative interactions, without changing the driving or coupling springs (“Configuration 1” in C). Graphs viivii-ixix require frustrated interactions; we obtain them with a second set of driving and coupling springs (“Configuration 2” in C). B. Switching thresholds calculated from the model, using the experimental parameters. C. Experimental parameters. We fix $x_{1}=0$ cm, $y=14.5$ cm, $L=6.3$ cm, $\ell=4.7$ cm, and $w=15.2$ cm. D. Schematic illustrating the role of interactions on hysteron 2 in graphs $i$-ixix. Top row: hysteron 1 is $(+)$; bottom row: hysteron 1 is ()(-). Smaller cooperative interactions are required to create graphs $i$-vivi in B, whereas larger frustrated interactions are needed for graphs viivii-ixix.
Figure S2: Diagram showing equality of $J_{ij}^{+}$ and $J_{ij}^{-}$ when $\theta_{i}^{+}=-\theta_{i}^{-}$ and $\ell_{i}=-\ell_{i}^{\prime}$ for all $i$. A. Top schematic: Generic configuration with torque $\tau_{12}(+,-)$ on rotor $1$. Locations of both parallel and crossed springs are illustrated, but only one set is present at a time. If $\theta_{i}^{+}=-\theta_{i}^{-}$ and $\ell_{i}=-\ell_{i}^{\prime}$ for $i=1,2$, then reflecting this schematic about the horizontal is equivalent to flipping the states of both rotors, implying $\tau_{12}(-,+)=-\tau_{12}(+,-)$. B. Analogous pair of schematics showing $\tau_{12}(-,-)=-\tau_{12}(+,+)$ under the same conditions. Using Eq. 3 in the main text, we have $J_{12}^{+}=(\tau_{12}(-,+)-\tau_{12}(-,-))/C^{-}=(\tau_{12}(+,+)-\tau_{12}(+,-))/C^{+}=J_{12}^{-}$, where $C^{\mp}=2Lk_{i}\cos\theta_{i}^{\mp}$.
Figure S3: Cooperative and frustrated interactions. A. We use the same system parameters as in Fig. 2 and calculate switching thresholds from the kinematic model (solid lines, Eq. 2) and the small-$\theta$ model (dashed lines, Eqs. 1, 6, and 7). The small-$\theta$ model captures the gross behavior of the system in spite of the large $\theta_{i}$ involved. B. Same analysis but with crossed springs, producing a frustrated interaction ($J_{ij}<0$) so that now $\gamma_{i}^{\pm}(+)>\gamma_{i}^{\pm}(-)$.